# Another look at inference after prediction

### v5

@@ Abstract
From structural biology to epidemiology, predictions from machine learning (ML) models increasingly complement costly gold-standard data to enable faster, more affordable, and scalable scientific inquiry. In response, prediction-based (PB) inference has emerged to accommodate statistical analysis using a large volume of predictions together with a small amount of gold-standard data. The goals of PB inference are two-fold: (i) to mitigate bias from errors in predictions and (ii) to improve efficiency relative to classical inference using only the gold-standard data. While early PB inference methods focused on bias, their ability to enhance efficiency remains a focus of ongoing research. We revisit a foundational PB inference method and show that a simple modification can be applied to guarantee provable improvements in efficiency. In doing so, we establish new connections between augmented inverse probability weighted estimators (AIPW) and several recently proposed PB inference methods with a similar focus. The utility of our proposal, which leverages prediction-based outcomes to enhance efficiency, is demonstrated through extensive simulation studies and an application to real data from the UK Biobank. Further, we contextualize PB inference by drawing connections to historical literature from economics and statistics, highlighting how classic methods directly inform this contemporary problem.
@@ 1 Introduction
Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to measure directly and then conducting statistical analyses which treat the predictions as observed data. 1, 2, 3, 4 This prediction-based (PB) inference approach, introduced by Wang et al,5 is comprised of two steps outlined in Figure 1. First, the outcome of interest is imputed using a prediction from a pre-trained model. Then, the predictions are used to draw inferences about the parameters of interest, such as the association parameters between the outcome and a set of covariates, together with a small amount of gold-standard data.***PB inference has primarily focused on outcomes that are difficult to measure. Recent work has considered predicted covariates (e.g., Kluger et al6). In contrast to classical inference that relies solely on gold-standard data, PB inference aims to leverage the predictions to improve statistical efficiency without introducing bias.
@@
In the last several years, applications of PB inference have arisen in diverse fields, spanning genetics, biology, and medicine. For example, population biobanks, such as the UK Biobank ($n\approx$ 500K individuals), collect high-throughput genomic data on large populations and offer unprecedented opportunities for genetic discovery. However, many phenotypes are only partially observed due to the time and expense of measurement. ML models are frequently used to impute missing outcomes based on fully observed baseline data to enable downstream association studies with phenotypes of interest. 7, 8, 9, 10, 11, 12, 13, 14 Similarly, in structural biology, AlphaFold provides protein conformation predictions without the need for expensive and labor-intensive experiments. 15, 16 In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review by trained annotators, facilitating clinical and epidemiological studies. 17, 18, 19
@@
Figure 1: Prediction-based (PB) inference. $Y$ is the outcome of interest, which is only partially observed, $\bf{Z}$} is a set of predictors of $Y$, $\widehat{Y}$ is the prediction of $Y$ based on $\bf{Z}$}, and $\bf{X}$} is a set of covariates. The focus of scientific inquiry is on the association between $Y$ and $\bf{X}$}. Classical statistical inference relies solely on the data on $Y$ and $\bf{X}$} while PB inference incorporates $\widehat{Y}$ to make use of all of the available data. Dark red represents labeled observations while light red represents unlabeled observations.
@@
Given the relevance of PB inference, Motwani and Witten 20 recently revisited two seminal methods in the field. In the context of least squares estimation, they showed that the post-prediction inference (PostPi) method proposed by Wang et al5 provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios (e.g., when the pre-trained ML model perfectly captures the underlying regression function). This is a significant practical limitation. PB inference methods should guarantee valid inference (i.e., hypothesis tests with controlled type I error and confidence intervals with nominal coverage rates) in the common scenario that the prediction model is imperfect. In contrast to PostPi, the prediction powered inference (PPI) method proposed by Angelopoulos et al 21 yields consistent estimation and valid inference regardless of the quality of the ML model. This finding is critical given that the PPI paper, published just over a year ago in the high-impact journal Science, has already been cited nearly 150 times and widely adopted in applications.
@@
In this paper, we further study the PPI estimator by assessing its statistical efficiency. Our contributions are three-fold. First, we make new contributions to the literature identifying inefficiencies of the PPI method. 22, 23, 24 We derive a set of results that characterize when the PPI estimator gains over classical estimation using labeled data only in the context of Z-estimation. Second, we show that a more efficient estimator can be obtained by adding a simple weight to the PPI estimator. This approach was proposed by Chen & Chen (CC) over 25 years ago in the context of double sampling designs 25 and its application within the PB inference setting is new. Third, we draw important connections between augmented inverse probability weighted (AIPW) estimators, our proposal, and several recently proposed methods. In doing so, we identify settings in which the CC estimator is theoretically guaranteed to improve over comparable approaches. Building on these findings, we also revisit literature from statistics and economics to identify work dating back to the 1960s that bears directly on current developments in PB inference. Overall, our contributions put classical statistical methods in conversation with this modern problem to inform future methodological developments and applications.
@@
We begin by formally introducing the PB inference problem, reviewing the contributions of Motwani and Witten20 and summarizing existing work in Section 2. We then study the asymptotic properties of the PPI estimator, showing that it can be outperformed by classical inference methods and introduce the CC estimator in Section 3. Our key finding is that the CC estimator is optimal with respect to statistical efficiency in the class of estimators to which the PPI estimator belongs. In Section 4, we discuss related literature and provide connections between the CC estimator, other PB inference methods, and AIPW estimators. We also develop a set of results that compare the efficiency of the CC estimator with several PB inference methods and identify scenarios where the CC estimator is theoretically optimal. We then evaluate our findings with extensive simulation studies in Section 5. In Section 6, we investigate the association between android fat mass and clinical and demographic covariates in the South Asian population in the UK Biobank to demonstrate the practical performance of these approaches. We close with additional discussion in Section 7. Overall, our theoretical and numerical results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it a preferred choice for use in practice.
@@ 2.1 Notation and Assumptions
We let $Y$ denote the outcome of interest and $\bf{Z}$} its predictors. $Y$ is only partially observed and a pre-trained ML model is available to predict it over the full data given $\bf{Z}$}, which is fully observed. The ML model $\hat{f}:\mathcal{Z}\rightarrow\mathcal{Y}$ maps the predictor space $\mathcal{Z}$ to the outcome space $\mathcal{Y}$ and produces predictions $\widehat{Y}=\hat{f}({\bf{Z}})$. We let ${\bf{X}}=(1,X_{1},\dots,X_{p})^{{\sf\scriptscriptstyle{T}}}$ denote fully observed covariates that are potentially associated with $Y$ and that are the focus of scientific inquiry. $\bf{X}$} may also be predictive of $Y$ and these covariates may or may not be included in $\bf{Z}$}. The available data therefore consists of a labeled dataset and an independent unlabeled dataset, respectively denoted as
@@
where $n=n_{\text{\tiny{lab}}}+n_{\text{\tiny{unlab}}}$, $n_{\text{\tiny{unlab}}}$ is the number of unlabeled observations and $n_{\text{\tiny{lab}}}/n\to\pi$ as $n\to\infty$ for $\pi\in(0,1)$.
We focus on the case in which $\mathscr{L}$ and $\mathscr{U}$ include independent and identically distributed samples from a common population of interest. Additionally, $\hat{f}$ is assumed to be pre-trained with a dataset that is independent of $\mathscr{L}$ and $\mathscr{U}$ and is therefore treated as “fixed” (i.e., the variation in estimating $\hat{f}$ is not of interest in downstream inference). $\hat{f}$ may be a black box function in that its training data and specification are not be available to the users. Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been considered.26
@@ 2.2 Parameter of interest
We focus on Z-estimation and aim to make inferences on the population parameter $\bm{\beta}^{*}$ that is the solution to the estimating equation
$\mathbb{E}[\bm{\phi}(Y,{\bf{X}};\bm{\beta})]=\bm{0}.$
$\bm{\phi}(\cdot)$ is a deterministic function that satisfies conditions C1-C4 in the Appendix. The simplest consistent estimator of $\bm{\beta}^{*}$, which we denote by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, uses only $\mathscr{L}$ for estimation and is the solution to the estimating equation
$\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta})=\bm{0}.$
@@ 2.3 Motivation for PB inference
Although $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is consistent and enables valid inference for $\bm{\beta}^{*}$, it has limited precision when gold-standard data is difficult to collect and $n_{\text{\tiny{lab}}}$ is not large. In contrast, the predictions $\widehat{Y}$ can be computed in the full data, leading to a much larger dataset for analysis. However, relying on $\widehat{Y}$ for inference may result in biased estimation depending on the quality of predictions from $\hat{f}$. For example, in “naive” inference, $\widehat{Y}$ is directly substituted for $Y$ in downstream analysis.5 The parameter $\bm{\gamma}^{*}$ targeted by this approach solves
$\mathbb{E}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma})]=\bm{0}.$
The corresponding estimator based on $\mathscr{L}$ and $\mathscr{U}$, denoted as $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, is the solution to
$\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}.$
As the distribution of $(\widehat{Y},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ may not coincide with that of $(Y,{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$, $\bm{\gamma}^{*}$ bears no necessary correspondence with $\bm{\beta}^{*}$ except under restrictive scenarios. For example, $\bm{\gamma}^{*}$ coincides with $\bm{\beta}^{*}$ when $\bf{Z}$} contains $\bf{X}$}, the ML model captures the true regression function $\mathbb{E}(Y\mid{\bf{Z}})$, and the focus of inference is on a parameter from a linear regression model. To overcome these limitations, PB inference methods utilize $\mathscr{L}$ and $\mathscr{U}$ to essentially balance the relative contributions of $Y$ and $\widehat{Y}$ in making inference on $\bm{\beta}^{*}$.
@@ 2.4 PB inference methods
In the context of least squares estimation, Motwani and Witten20 revisited two seminal strategies for PB inference. One of these is the PostPI method of Wang et al,5 which was the first modern method proposed for PB inference. PostPI consists of two steps to incorporate $\widehat{Y}$ into inference for $\bm{\beta}^{*}$. First, $\mathscr{L}$ is used to model the relationship between $Y$ and $\widehat{Y}$ to address potential inaccuracies in the predictions. Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$ with either analytical or bootstrap methods. This approach was shown to yield a consistent estimator for $\bm{\beta}^{*}$ under the strong and arguably unrealistic assumption that the ML model captures the true regression function. When this assumption is not met, the PostPI estimator may not be consistent. Consequently, corresponding hypothesis tests may not have controlled type I error rates and confidence intervals may not have adequate coverage.
Motwani and Witten20 also studied the PPI method 21 which was proposed a few years after PostPI. Unlike its predecessor, PPI has the crucial property that inference remains valid regardless of the quality of the predictions. That is, $\hat{f}$ need not capture the true regression function for the estimator to maintain consistency for $\bm{\beta}^{*}$. The PPI estimator utilizes a common statistical technique that augments an initial consistent estimator for the target parameter based solely on $\mathscr{L}$ with an additional term that is estimable from both $\mathscr{L}$ and $\mathscr{U}$. The augmentation term converges in probability to zero in order to improve efficiency without introducing bias. This approach dates back to at least Robins et al 27 and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem. 25, 28, 29, 30, 31 We return to this point in our review of related literature in Section 4.
Formally, the PPI estimator is
@@
where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ solves $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$ and $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ solves $\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$Angelopoulos et al21 express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$. We reframe the PPI method to study its statistical efficiency.. In contrast to the PostPI method, Motwani and Witten20 show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is consistent for $\bm{\beta}^{*}$ and $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*})$ converges weakly to a zero-mean multivariate normal distribution regardless of the quality of the prediction. Analogous results for Z-estimation problems are established in the original PPI paper (e.g., see Theorem S1 in the Supplementary Materials). This approach therefore asymptotically controls the type I error rate in hypothesis tests and produces confidence intervals with the nominal coverage level.
However, for a PB inference method to be preferred in practice over classical inference based on $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, it should not only yield valid inference, but also be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\widehat{Y}$ is informative of $Y$ and equally efficient when $\widehat{Y}$ is uninformative of $Y$. Two recent works have empirically identified that the PPI method can be less efficient than classical inference and have proposed extensions of the original method that are theoretically guaranteed to outperform classical inference. The first such extension, called PPI++, abandons the augmentation-based approach of PPI and instead directly augments the estimating equation for $\bm{\beta}^{*}$ with information from $\widehat{Y}$. A tuning parameter is used that controls the amount of information from $\widehat{Y}$ that is used for estimation and is cleverly designed to minimize the trace of the resulting estimator’s asymptotic covariance matrix. Subsequently, Miao et al. 23 proposed Post-Prediction Adaptive inference (PSPA), which adopts a similar strategy, but instead produces an estimator with minimal element-wise variance. While both PPI++ and PSPA offer provable improvements over classical inference using only $\mathscr{L}$, their relationships to the original PPI framework remain underexplored. Furthermore, the conditions under which these methods are optimal have yet to be fully established. We revisit these proposals and clarify some of these connections in Section 4.

### v6

@@ Abstract
From structural biology to epidemiology, predictions from machine learning (ML) models are increasingly used to complement costly gold-standard data to enable faster, more affordable, and scalable scientific inquiry. In response, prediction-based (PB) inference has emerged to accommodate statistical analysis using a large volume of predictions together with a small amount of gold-standard data. The goals of PB inference are two-fold: (i) to mitigate bias from errors in predictions and (ii) to improve efficiency relative to traditional inference using only the gold-standard data. While early PB inference methods focused on bias, their ability to enhance efficiency remains unclear. We revisit a popular PB inference method and show that a simple modification can be applied to guarantee improvements in efficiency beyond yielding valid inferences when the ML predictions are imperfect. The utility of this approach in leveraging prediction-based outcomes to enhance efficiency is demonstrated through extensive simulation studies and an application to the UK Biobank data. We further contextualize the problem of PB inference through historical literature from economics and statistics to highlight perspectives from classical methods in this contemporary problem.
@@ 1 Introduction
Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to measure directly and conducting statistical analyses treating the predictions as observed data. hoffman2024we, fan2024narratives, gao2024semi, miao2024task This prediction-based (PB) inference approach, introduced by Wang et al,wang2020methods is comprised of two steps outlined in Figure 1. First, the outcome of interest is imputed using a prediction from a pre-trained model. Then, the predictions are used to draw inferences about the parameters of interest, such as the association parameters between the outcome and a set of covariates, together with a small amount of gold-standard data.***PB inference has primarily focused on outcomes that are difficult to measure. Recent work has considered predicted covariates (e.g., Kluger et alkluger2025prediction). In contrast to traditional inference that relies solely on gold-standard data, PB inference aims to leverage the predictions to improve statistical efficiency without introducing bias.
@@
In the last several years, applications of PB inference have arisen in diverse fields, spanning genetics, biology, and medicine. For example, population biobanks, such as the UK Biobank ($n\approx$ 500K individuals), collect high-throughput genomic data on large populations and offer unprecedented opportunities for genetic discovery research. However, many traits and disease outcomes are only partially observed due to the time and cost of measurement. ML models are commonly used to impute missing outcomes based on fully observed baseline data to enable downstream association studies. tong2020augmented, alipanahi2021glaucoma, cosentino2023copd, dahl2023phenotype, an2023autocomplete, mccaw2024synthetic, liu2024semiparametric, miao2024popgwas Similarly, in structural biology, AlphaFold provides structural predictions without the need for expensive and labor-intensive experiments for proteomic studies. jumper2021highly, gan2022activation In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review for epidemiological and clinical studies. yang2022large, zhou2022cancerbert, alsentzer2023zero
@@
Figure 1: Prediction-based (PB) inference. $Y$ is the outcome of interest, which is only partially observed, $\bf{Z}$} is a set of predictors of $Y$, $\widehat{Y}$ is the prediction of $Y$ based on $\bf{Z}$}, and $\bf{X}$} is a set of covariates. The focus of scientific inquiry is on the association between $Y$ and $\bf{X}$}. Traditional statistical inference relies solely on the data on $Y$ and $\bf{X}$} while PB inference incorporates $\widehat{Y}$ to make use of all of the available data. Dark red represents labeled observations while light red represents unlabeled observations.
@@
Given the relevance of PB inference, Motwani and Witten motwani2023revisiting recently revisited two seminal methods in the field. In the context of least squares estimation, they showed that the post-prediction inference (PostPi) method proposed by Wang et alwang2020methods provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios (e.g., when the pre-trained ML model perfectly captures the underlying regression function). This is a significant practical limitation as it is essential that PB inference methods yield valid inference (i.e., hypothesis tests with controlled type I error and confidence intervals with nominal coverage rates) in the likely scenario that the prediction model is imperfect. In contrast, the prediction powered inference (PPI) method proposed by Angelopoulos et al angelopoulos2023prediction yields consistent estimation and valid inference regardless of the quality of the ML model. This finding is critical given that the PPI paper, published just over a year ago in the high-impact journal Science, has already been cited nearly 150 times and widely adopted in applications.
@@
In this paper, we build on the work of Motwani and Witten motwani2023revisiting and further examine the PPI estimator by assessing its statistical efficiency. Our contributions are two-fold. First, we develop a set of results that characterize when the PPI estimator gains over traditional estimators using labeled data only. We also show that a more efficient estimator can be obtained by adding a simple weight to the PPI estimator. This approach was proposed by Chen & Chen (CC) over 25 years ago in the context of double sampling designs.chen2000unified Second, we revisit literature from statistics and economics and identify work dating back to the 1960s that bears directly on current developments in PB inference. Our review puts classical statistical methods in conversation with this modern problem to inform future methodological advances and applications.
@@
We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Wittenmotwani2023revisiting in Section 2. We then study the asymptotic properties of the PPI estimator, showing that it can be outperformed by traditional inference methods, and introduce the CC estimator in Section 3. Our key finding is that the CC estimator is optimal with respect to statistical efficiency in a class of estimators to which the PPI estimator belongs. In Section 4, we discuss related literature and provide connections between the CC estimator and other established methods. We then evaluate our findings with extensive simulation studies in Section 5. In Section 6, we investigate the association between android fat mass and clinical and demographic covariates in the South Asian population in the UK Biobank to demonstrate the practical performance of these approaches. We close with additional discussion in Section 7. Overall, our theoretical and numerical results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it a preferred choice for use in practice.
@@ 2.1 Notation and Assumptions
We let $Y$ denote the outcome of interest and $\bf{Z}$} a set of predictors. $Y$ is only partially observed and a pre-trained ML model is available to predict it over the full data given $\bf{Z}$}, which are fully observed. The ML model $\hat{f}:\mathcal{Z}\rightarrow\mathcal{Y}$ maps the predictor space $\mathcal{Z}$ to the outcome space $\mathcal{Y}$ and produces predictions $\widehat{Y}=\hat{f}({\bf{Z}})$. We let ${\bf{X}}=(1,X_{1},\dots,X_{p})^{{\sf\scriptscriptstyle{T}}}$ denote fully observed covariates that are potentially associated with $Y$ and that are the focus of scientific inquiry. $\bf{X}$} may also be predictive of $Y$ and these covariates may or may not be included in $\bf{Z}$}. The available data therefore consists of a labeled dataset and an independent unlabeled dataset, respectively denoted as
@@
where $n=n_{\text{\tiny{lab}}}+n_{\text{\tiny{unlab}}}$, $n_{\text{\tiny{unlab}}}$ is the number of unlabeled observations, and $n_{\text{\tiny{lab}}}/n\to\pi$ as $n\to\infty$ for $\pi\in(0,1)$.
We focus on the case in which $\mathscr{L}$ and $\mathscr{U}$ include independent and identically distributed samples from a common population of interest. Additionally, $\hat{f}$ is assumed to be pre-trained with a dataset that is independent of $\mathscr{L}$ and $\mathscr{U}$ and is therefore treated as “fixed” (i.e., the variation in estimating $\hat{f}$ is not of interest in downstream inference). $\hat{f}$ may be a black box function in that its training data and specification are not be available to the users. Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been considered.zrnic2024cross
@@ 2.2 Parameter of interest
Following Motwani and Witten,motwani2023revisiting we consider least squares estimation and aim to make inferences on the population parameter
$\bm{\beta}^{*}\coloneq\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}[(Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta})^{2}]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}({\bf{X}}Y),$
which quantifies the association between $Y$ and $\bf{X}$}. The simplest consistent estimator of $\bm{\beta}^{*}$, which we denote by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, uses only $\mathscr{L}$ for estimation and solves the estimating equation
$\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta})\coloneq\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\beta})=\bm{0}.$
@@ 2.3 Motivation for PB inference
Although $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is consistent and enables valid inference for $\bm{\beta}^{*}$, it has limited precision when gold-standard data is difficult to collect and $n_{\text{\tiny{lab}}}$ is not large. In contrast, the predictions $\widehat{Y}$ can be computed in the full data, leading to a much larger dataset for analysis. However, relying on $\widehat{Y}$ for inference may result in biased estimation depending on the quality of predictions from $\hat{f}$. For example, in “naive” inference, $\widehat{Y}$ is directly substituted for $Y$ in downstream analysis.wang2020methods The parameter targeted by this approach is
[…]
and the corresponding least squares estimator, denoted as $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, is the solution to
$\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})\coloneq\sum_{i=1}^{n}{\bf{X}}_{i}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})=\bm{0}.$
As the distribution of $(\widehat{Y},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ may not coincide with that of $(Y,{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$, $\bm{\gamma}^{*}$ bears no necessary correspondence with $\bm{\beta}^{*}$ except under restrictive scenarios (e.g., when $\bf{Z}$} contains $\bf{X}$} and the ML model captures the true regression function). To overcome these limitations, PB inference methods utilize $\mathscr{L}$ and $\mathscr{U}$ to essentially balance the relative contributions of $Y$ and $\widehat{Y}$ in making inference on $\bm{\beta}^{*}$.
@@ 2.4 PB inference methods
Motwani and Wittenmotwani2023revisiting revisited two seminal strategies for PB inference. The first is the PostPI method of Wang et al,wang2020methods which is a two-step procedure to incorporate $\widehat{Y}$ into estimation. First, $\mathscr{L}$ is used to model the relationship between $Y$ and $\widehat{Y}$ to address potential inaccuracies in the predictions. Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$ with either analytical or bootstrap approaches. This approach is shown to yield a consistent estimator for $\bm{\beta}^{*}$ under the strong and unrealistic assumption that the ML model captures the true regression function. When this assumption is not met, the PostPI estimator may not be consistent. Consequently, corresponding hypothesis tests may not have controlled type I error rates and confidence intervals may not have adequate coverage.
Motwani and Wittenmotwani2023revisiting also considered the PPI method,angelopoulos2023prediction which has the crucial property that inference remains valid regardless of the quality of the predictions. That is, the PPI method does not require that $\hat{f}$ capture the true regression function for the estimator to maintain consistency for $\bm{\beta}^{*}$. The proposed estimator utilizes a common statistical technique involving the augmentation of an initial consistent estimator for the parameter of interest with an additional term that converges in probability to 0. This approach dates back to at least Robins et al robins1994estimation and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem. chen2000unified, tsiatis2006, tian2012covariate, wang2015semiparametric We return to this point in our review of related literature in Section 4.
The PPI estimator is
@@
where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ solves $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$ and $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ solves $\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$Angelopoulos et alangelopoulos2023prediction express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$. We reframe the PPI method to study its statistical efficiency.. In contrast to the PostPI method, Motwani and Wittenmotwani2023revisiting show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is consistent for $\bm{\beta}^{*}$ and $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*})$ converges weakly to a zero-mean multivariate normal distribution regardless of the quality of the prediction. This approach therefore asymptotically controls the type I error rate in hypothesis tests and produces confidence intervals with the nominal coverage level.